OpenStudy (anonymous):

whats the correct way to write 55% as a fraction

OpenStudy (anonymous):

0.55

OpenStudy (anonymous):

55 over 100

OpenStudy (goformit100):



OpenStudy (anonymous):

55/100

OpenStudy (anonymous):

55%=55/100 55%=0.55

OpenStudy (unklerhaukus):

\[\frac{55\%}{1}\]

OpenStudy (unklerhaukus):

\[\uparrow\]incorrect way

OpenStudy (anonymous):

55*100/100*100